I need help with this problem. I am not sure what formula to use.
An aquarium tank can hold 540 liters of water. There are two pipes that can be used to fill the tank. The first pipe alone can fill the tank in 60 minutes. The second pipe can fill the tank in 90 minutes by itself. When both pipes are working together, how long does it take them to fill the tank?

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(57278)

An aquarium tank can hold 540 liters of water.
There are two pipes that can be used to fill the tank. The first pipe alone can fill the tank in 60 minutes.
The second pipe can fill the tank in 90 minutes by itself.
When both pipes are working together, how long does it take them to fill the tank?
-------------
1st pipe DATA:
Time = 60 minutes/job ; rate = 1/60 job/min
-------------------------------------
2nd pipe DATA:
Time = 90 minutes/job ; rate = 1/90 job/min
-------------------------------------
Together DATA:
Time = x minutes/job ; rate = 1/x job/min
----------------------------------------
EQUATION:
rate + rate = together rate
1/60 + 1/90 = 1/x
(90+60)/(60*90) = 1/x
x = 5400/150
x = 36 minutes (time to do the job together)
======================================
Cheers,
Stan H.

Answer by Alan3354(30993)

I need help with this problem. I am not sure what formula to use.
An aquarium tank can hold 540 liters of water. There are two pipes that can be used to fill the tank. The first pipe alone can fill the tank in 60 minutes. The second pipe can fill the tank in 90 minutes by itself. When both pipes are working together, how long does it take them to fill the tank?
-------------------------
There are 2 ways to work this. One way depends on knowing the volume of the tank, the 540 liters. The 2nd way is more general, and the volume value isn't needed.
--------------
Method 1
540 liters in 60 mins = 9 liters/minute
540 liters in 90 mins = 6 liters/minute
Total = 15 liters/min
540/15 = 36 minutes.
-------------
Method 2
Pipe 1 does the volume V in 60 mins, or V/60 per minute.
Pipe 2 does V/90 per minute.
Together, they do V/60 + V/90
= 3V/180 + 2V/180 = 5V/180 per minute
The time is the inverse, 180/5 = 36 minutes. Same answer, but independent of the volume.
---------
There's also a shortcut when there are 2 pipes: The product divided by the sum.
60*90/(60+90) = 5400/150 = 36. That can save time on a test.